OpenStudy (aroub):

I just want to check my answer: When traveling with the wind, a plane flies 300 miles between Chicago and Columbus in 40 minutes. It returns against the wind in 45 minutes. Find the air speed of the plane and the rate of the wind. air speed of the plane: 85/12 and rate of the wind: 5/12 miles/min

OpenStudy (experimentx):

(x+y)*40 = 300 (x-y)*45 = 300 http://www.wolframalpha.com/input/?i=%28x%2By%29*40+%3D+300%3B+%28x-y%29*45+%3D+300

OpenStudy (anonymous):

In math classes, we sometimes measure aircraft speed as miles/minute, but out in real life, everything is measured in miles per hour. That means 40 minutes is 2/3 hour, and 45 min is 3/4 hour.

OpenStudy (phi):

Your answers are correct, but generally people use mph (miles per hour), so 425 mph and 400 mph

OpenStudy (aroub):

Oh, okay. Thank youuu =D